# MIMIR: Masked Image Modeling for Mutual Information-based Adversarial Robustness

### v2

@@ 3.3 Design Details
where $\theta$ are the parameters of the autoencoder.
Adversarial fine-tuning. After the autoencoder is trained, we discard the decoder and initialize a classification layer for the encoder to build a complete model. We use simple adversarial training, PGD [55], to fine-tune our pre-trained models so that we can ensure the improvements come from the MIMIR pre-training. The complete model is trained at the fine-tuning stage.
@@ 3.4 Theoretical Justification
In this section, we provide theoretical justification showing that the MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained. Let $F$ be any classifier trained on clean samples with desirable prediction accuracy, which may suffer from adversarial attacks. We begin our analysis by first presenting Lemma 3.1.
Lemma 3.1
Given an arbitrary classifier $F$, let $F(x+\delta)$ and $F(x_{re})$ denote, respectively, the predicted labels of adversarial sample $x+\delta$ and reconstructed sample $x_{re}$. We have:
@@
which indicate that $F(x+\delta)$ is an indirect observation of $x+\delta$, whereas both $F(x_{re})$ and $x_{re}$ are indirect observations of $z$.
@@
Now, we define $p_{e}$ as the probability that the predicted label of $x+\delta$ by $F$ is not equal to that of $x_{re}$, i.e., $p_{e}=\mathbb{P}(F(x+\delta)\neq F(x_{re}))$. Intuitively, our autoencoder is trained to only recover clean sample $x$ without any interference from $\delta$. Hence, a relatively large value of $p_{e}$ is expected. In the following, we establish the connection between $p_{e}$ and $I(x+\delta,z)$ with both lower and upper bounds, showing that minimizing $I(x+\delta,z)$ also encourages a large value of $p_{e}$.
Proposition 3.2
@@
where $|F(x+\delta)|$ is the total number of categories [^3] For instance, for CIFAR-10 [46], $|F(x+\delta)|=10$..
Proof:
@@
By applying Fano’s inequality [30, 58], we obtain:
@@
Therefore, we obtain a lower bound of $I(x+\delta,z)$. If we use CIFAR-10 ($|F(x+\delta)|=10$) and assume the predicted labels $F(x+\delta)$ follow a uniform distribution, we can visualize the lower bound as a function of $p_{e}$ as shown in Figure 2, from which we observe an obvious monotonic inverse relationship between $I(x+\delta)$ and $p_{e}$ in the range $p_{e}\in[0,0.9]$.
In fact, we can also obtain an upper bound, under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq. (11).
Proposition 3.3
@@
Proof:
By the Hellman-Raviv inequality [38, 13], we have:
@@
Similar to the lower bound, the upper bound also indicates $I(x+\delta,z)$ is inversely proportional to $p_{e}$ as shown in Figure 2.
In fact, apart from the above-mentioned lower and upper bounds, there exists an alternative and intuitive way to understand the mechanism of minimizing $I(x+\delta,z)$. For simplicity, let us assume the clean data $x$ and adversarial perturbation $\delta$ are independent, then:
@@
According to [78], minimizing the expected reconstruction error between clean sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input. Therefore, by minimizing $I(x+\delta,z)$, the network is forced to minimize $I(\delta,z)$ (since $I(x,z)$ is maximized). In other words, only the information about $\delta$ has been removed from $z$. This also explains the robustness of $z$.
TABLE I: Different learning rates. Fine-tuned for 50 epochs.
@@
CIFAR-10 ViT-T 5.0e-4 76.30 47.60
1.0e-3 80.69 49.56
1.0e-2 85.62 48.78
5.0e-2 85.12 50.30
1.0e-1 84.51 50.40
TABLE II: Top 1 natural and adversarial accuracy on CIFAR-10 test set. The results of ARD+PRM [57] are taken from the original paper. As there is no ViT-T architecture, we use the same architecture as DeiT-T.
@@
CIFAR-10 [46] ($l_{\infty},\epsilon=8/255$) ViT-T (DeiT-T) ImageNet-1K PGD 75.46 48.10 47.96 43.62
MAE [36] PGD 79.89 48.64 48.43 44.27
ImageNet-1K ARD+PRM [57] 79.60 50.33 50.15 45.99
MIMIR PGD 84.82 57.51 57.38 52.96
ViT-S ImageNet-1K PGD 79.59 50.86 50.73 46.37
MAE [36] PGD 85.97 54.19 53.94 50.49
ImageNet-1K ARD+PRM [57] 81.86 51.73 51.46 47.33
MIMIR PGD 88.11 56.63 56.34 53.18
ViT-B ImageNet-1K PGD 83.16 52.98 52.71 49.06
MAE [36] PGD 88.86 57.03 56.86 52.42
ImageNet-1K ARD+PRM [57] 84.90 53.80 53.51 50.03
MIMIR PGD 89.30 58.07 57.87 54.55
ConViT-T ImageNet-1K PGD 53.09 33.63 33.61 29.65
MAE [36] PGD 75.49 45.29 45.17 41.12
ImageNet-1K ARD+PRM [57] 80.28 49.86 49.55 45.42
MIMIR PGD 80.74 49.37 49.16 45.04
ConViT-S ImageNet-1K PGD 54.03 34.61 34.60 30.60
MAE [36] PGD 86.43 54.76 54.59 51.13
ImageNet-1K ARD+PRM [57] 84.32 53.10 52.81 48.85
MIMIR PGD 87.49 56.35 56.20 52.54
ConViT-B ImageNet-1K PGD 61.54 38.77 38.71 34.21
MAE [36] PGD 89.03 58.19 57.75 54.25
ImageNet-1K ARD+PRM [57] 85.80 53.36 53.05 49.33
MIMIR PGD 89.30 59.31 59.01 55.64
4 Experimental Evaluation
4.1 Experimental Setup
We use PyTorch [60] 2.1.0 for implementation, and mixed precision [56] to accelerate training. We use three datasets to comprehensively evaluate MIMIR: CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26], details of datasets are given in Appendix A. Appendix B provides details about decoder hyperparameters.
Model Architecture: We evaluate MIMIR with three commonly used architectures with multiple scales: ViT [28], ConViT [24], and XCiT [3].
Training hyperparameters. For all experiments, we train models from scratch (unless otherwise specified). Following MAE [36], we do pre-training by MIMIR for 800 epochs and then fine-tuning for 100 epochs. The number of warmup epochs is 40 and 10 for pre-training and fine-tuning, respectively. We use AdamW [53] as an optimizer for both pre-training and fine-tuning. We use cosine decay as the learning rate scheduler.
At pre-training, one-step FGSM with random initialization achieves better performance than using ten-step PGD in Table VII. Therefore, we use a one-step algorithm to generate adversarial examples for all three datasets. The perturbation budget is $\epsilon=8,\alpha=10$. At fine-tuning stage, we use a 10-step PGD algorithm to generate adversarial examples for training to be aligned with baselines. The perturbation limit is $\epsilon=8,\alpha=2$. When fine-tuning with ImageNet-1K, we use one-step FGSM with random initialization to align with baselines and for better efficiency. Following previous research [86, 87], we fine-tuned twice for two different perturbation limits, $\epsilon=2,\epsilon=4$.
As discussed in Section 4.3, strong data augmentation is harmful to adversarial training at a common training schedule (50 or 100 epochs for fine-tuning). We only use simple augmentations, including random resized crops and random horizontal flips.
Previous research used rather different fine-tuning learning rates, so we chose the learning rate by testing different values as given in Table I. We use the best, i.e., $25.0e-$ 2, as the base learning rate (blr) for CIFAR-10. The learning rate is calculated based on blr, i.e., $lr=blr\times\frac{batchsize}{256}$. For Tiny-ImageNet and ImageNet-1K, we use the same blr as MAE [36], i.e., $0.001$. At the pre-training stage, we use $41.5e-$ 4 as the learning rate for all datasets. In addition, we noticed that it is possible to get NaN (Not a Number) for loss in a few cases while adversarial training, i.e., the training fails. We solve this problem by using a learning rate that is ten times smaller. More details about hyperparameters can be found in the Appendix D.
TABLE III: Top 1 natural and adversarial accuracy on ImageNet-1K.
Dataset Models Pre-train Fine-tune $\epsilon=2/255$ $\epsilon=4/255$
[…]
ImageNet-1K [26] ($l_{\infty}$) ViT-S - FastAT [86] 71.37 49.49 49.41 66.92 34.07 33.82
- AGAT [87] 70.62 49.00 48.85 66.10 33.62 33.40
MIMIR FastAT [86] 74.60 54.56 54.55 71.29 40.98 40.63
ViT-B - FastAT [86] 70.31 50.55 50.06 65.18 33.59 33.39
- AGAT [87] 70.41 51.23 51.11 67.93 34.94 34.78
MIMIR FastAT [86] 75.88 55.42 55.36 73.22 41.26 40.46
CaiT-XXS24 - FastAT [86] 72.84 54.31 54.26 68.77 32.14 32.09
- AGAT [87] 71.15 54.17 54.08 68.22 31.46 31.39
MIMIR FastAT [86] 73.39 53.39 53.37 69.90 40.53 40.22
CaiT-S36 - FastAT [86] 72.51 53.12 52.76 71.20 33.02 32.84
- AGAT [87] 72.69 53.66 53.48 71.06 33.46 33.17
MIMIR FastAT [86] 76.05 56.78 56.75 73.57 40.03 39.16
TABLE IV: Comparison of different architectures under 100-step PGD attack.
Dataset Models $\epsilon=2/255$ $\epsilon=4/255$
Natural PGD Natural PGD
ImageNet-1K ($l_{\infty}$) ViT-S 74.60 54.56 71.29 40.63
ConViT-S 74.72 55.65 72.08 42.62
XCiT-S 74.39 55.21 71.77 40.97
XCiT-M 73.63 53.04 72.01 42.39
[…]
Tiny-ImageNet [47] ($l_{\infty},\epsilon=8/255$) WRN-28-10 Gowal et al. [33] $\times$ 512 400 51.56 N/A N/A 21.56
WRN-28-10 Wang et at. [83] 1m 512 400 53.62 N/A N/A 23.40
[…]
ViT-S MIMIR 1m 512/256 800/100 63.33 26.37 26.17 22.52
ViT-B MIMIR 1m 512/256 800/100 65.06 25.41 25.03 22.82
ConViT-T MIMIR 1m 512/256 800/100 52.61 24.61 24.58 19.42
ConViT-S MIMIR 1m 512/256 800/100 62.29 26.39 26.30 22.88
Evaluation metrics. We use natural accuracy and adversarial accuracy as evaluation metrics. Natural accuracy refers to the accuracy of a model on clean and unmodified inputs. The adversarial accuracy measures the accuracy of a model under a specific adversarial attack. To show that the improvement is not from obfuscated gradient [5], we use PGD [55] and AutoAttack [20] to evaluate the robustness of trained models. PGD is one of the most popular adversarial attacks used for generating adversarial examples. We evaluate MIMIR under 20 and 100 steps of PGD attack. AutoAttack is an ensemble of diverse parameter-free attacks, including white-box and black-box attacks. In our experiments, we use standard version of AutoAttack that contains four attacks, including APGD-ce [20], APGD-t [20], FAB-t [19], and Square [4].
We use the same perturbation budget for our experiments to compare with previous works. We apply the commonly-used $l_{\infty}$ norm for PGD [55] and AutoAttack (AA) [20]. We set the perturbation radius $\epsilon$ to $8/255$ for CIFAR-10 [46] and Tiny-ImageNet [47]. The step length is $\alpha=2/255$. As the image size of ImageNet-1K is much larger, we use different scales of $\epsilon$ for adversarial training and evaluation [26]. Following previous works [57, 87, 50], we use two perturbation radius $2/255,4/255$ for evaluation. The step length is $\alpha=1/255$.
4.2 Experimental Results
Overall performance. We apply MIMIR to CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26]. The results are shown in Tables IIIII, and V, respectively. Overall, MIMIR significantly improves adversarial accuracy under PGD and AA attacks. While training with CIFAR-10 on ViTs (ViT-S and ViT-B), MIMIR improves the accuracy by an average of 5.50% compared to the baseline. On ConViTs (ConViT-T, ConViT-S, and ConViT-B), MIMIR improves the accuracy by an average of 2.87% compared to the baseline. While using ImageNet-1K on ViTs (ViT-S and ViT-B), MIMIR improves the accuracy by an average of 5.52% compared to the baseline. When fine-tuning with clean data, MIMIR also shows improved natural accuracy in Table XIV. In addition, we also evaluate the time consumption and memory usage of MIMIR in Table X. We compare MIMIR with FastAT [86] and 10-step PGD adversarial training. The results show that using MIMIR for pre-training is much more efficient than traditional adversarial training.
[…]
Comparison with related work. We present the performances of MIMIR compared with recent baseline methods in Tables II and III. In Table II, we perform experiments using CIFAR-10 [46]. The results are compared with three baselines, including pre-training on ImageNet-1K and fine-tuning with PGD, MAE [36], and ARD+PRM [57]. The “ImageNet-1K” under the Pre-train column refers to using models pre-trained on ImageNet-1K, then fine-tuned with PGD for 40 epochs, which is the first baseline in ARD+PRM [57]. The MAE [36] baseline is pre-trained using clean CIFAR-10 images and fine-tuned with 10-step PGD adversarial training. The training schedule of MAE [36] baseline is 800 epochs for pre-training and 100 epochs for fine-tuning. Our MIMIR uses the same training schedule as MAE [36]. We provide results on fine-tuning with clean images in Appendix E. The ARD+PRM [57] baseline refers to the best algorithm in [57], where we use the results from that work. Note that ARD+PRM [57] baseline uses models pre-trained on ImageNet-1K. Using models pre-trained on ImageNet-1K and then fine-tuning for CIFAR-10 can easily get accuracy higher than 95% [75].
As MIMIR only uses CIFAR-10 for training, it is more challenging to achieve good performance. For a more fair comparison, we use elucidating diffusion model (EDM) data as a data augmentation since the baselines also use EDM data (Wang et at. [83] and Rebuffi et al. [62]) or models pre-trained on ImageNet-1K (ARD+PRM [57]). Generative data is usually used to improve adversarial training [83, 61, 63, 33]. Specifically, we use 5 million generated CIFAR-10 data and 1 million Tiny-ImageNet data provided by [83]. Note that the EDM data is generated only based on CIFAR-10 or Tiny-ImageNet, and the MAE baseline uses EDM data as well. The EDM data is applied to experiments with CIFAR-10 and Tiny-ImageNet but not to ImageNet-1K.
According to results in Table II, MIMIR consistently improves adversarial robustness compared to all baselines. It also improves the natural accuracy compared to the MAE baseline. This is because self-supervised MIMIR does not suffer from inconsistency between the ground truth and adversarial example [59]. Let us consider a joint data distribution $p_{d}(x,y)$ (training data) and a discriminative model $p_{\theta}(y|x)$, where $x$ is the data with its label $y$. In standard training, we hope the model learns discriminative features by minimizing the distance between $p_{d}(y|x)$ and $p_{\theta}(y|x)$. In adversarial training, we hope the model is robust by minimizing the distance between $p_{d}(y|x)$ and $p_{\theta}(y|x+\delta)$, i.e., Eq. (3). This means $p_{\theta}(y|x+\delta)$, instead of $p_{\theta}(y|x)$, is encouraged to converge to $p_{d}(y|x)$. Supervised adversarial training builds a shortcut from $x+\delta$ to $y$. For the optimal $\theta^{*}$ of minimizing Eq. (3), $p_{\theta^{*}}(y|x)\neq p_{d}(y|x)$ [59]. Therefore, there is always a trade-off between natural and adversarial accuracy in supervised adversarial training. However, such inconsistency does not apply to self-supervised MIMIR, as there is no label for self-supervised training. MIMIR encourages $p_{\theta_{ae}}(x_{re}|x+\delta)$ to converge to $p_{d}(x)$, which does no affect convergence of $p_{\theta_{ae}}(x_{re}|x)$ to converge to $p_{d}(x)$. This implies that supervised pre-training is harmful to natural accuracy, which is also aligned with the findings of previous research [39].
Table III shows the results evaluated on the ImageNet-1K [26] dataset. The performance of MIMIR is compared with two baselines, including FastAT [86] and AGAT [87]. The FastAT [86] is designed for efficient adversarial training so that it can be applied to large models and large datasets, such as ImageNet-1K. The main idea of FastAT is to use a single-step adversarial attack to generate adversarial examples. The authors refined the standard FGSM [32] by applying a random start at the beginning of the algorithm and a much larger step size to reach the perturbation limit. According to empirical results, they used $\alpha=1.25\times\epsilon$ for step size, which provided comparable performance to the best-reported result from free adversarial training [25]. The results of FastAT [86] and AGAT [87] are taken from [87]. The baselines of FastAT [86] and AGAT [87] are trained for 300 epochs with a cosine decay learning rate schedule, and the first 20 epochs for learning rate warmup. As shown in Table III, the accuracy of AGAT [87] may decrease compared to FastAT [86]. This is because AGAT [87] discards low-attention image patches for efficiency. The image patches, even with the low-attention scores, are useful for classification. In Table IV, we compare the performance among three architectures with multiple scales. MIMIR shows similar performance on other architectures as on the original ViTs, which means MIMIR is well-generalized to different architectures.
Figure 3: The loss and natural accuracy for three different augmentation strategies, using CIFAR-10 and ViT-small.
As CIFAR-10 is small compared to ImageNet-1K, we also provide experiments on Tiny-ImageNet for completeness. We use the same training schedule and hyperparameters for experiments on Tiny-ImageNet. As AGAT [87] and ARD+PRM [57] do not provide experiments on Tiny-ImageNet, we compare our results with the most recent adversarial robustness results from Wang et at. [83] and Rebuffi et al. [62]. They use generated data as augmentation for more robust adversarial training. Specifically, Gowal et al. [33] study four methods for generated data as augmentation, including Denoising Diffusion Probabilistic Model (DDPM) [41], Very Deep Variational Auto-Encoder (VDVAE) [18], StyleGAN [45], and BigGAN [12]. Wang et at. [83] demonstrated better diffusion models and further improved adversarial robustness compared to previous works. In Table V, we show our experimental results compared with previous works.
Our method provides competitive performance. Especially when the baselines only use generated data based on Tiny-ImageNet, MIMIR provides the best natural accuracy (65.06%). When the baselines use generated data based on ImageNet-1K, and we only use that based on Tiny-ImageNet, MIMIR still shows comparable performance. In addition, according to Table X, the MIMIR training schedule (800 pre-training + 100 fine-tuning) takes approximately 200 fine-tuning (10-step PGD) epochs of training time. Thus, MIMIR costs much less training time and gets comparable results to baselines.
TABLE VI: Data augmentation ablation with longer fine-tuning schedule.
[…]
ViT-B 800 Weak Augmentation 89.90 60.26
+ CutMix [95],MixUp [97] 91.01 60.62
+ Randaugment [21] 90.19 62.75
4.3 Data Augmentation is Harmful
As mentioned in previous works [6, 25], strong data augmentation makes the training samples too difficult to be learned by ViTs while conducting adversarial training. The strong data augmentation refers to the combination of Randaugment [21], CutMix [95], and MixUp [97]. In this section, we evaluate two different solutions to ease this problem. First, we only use simple data augmentation for adversarial training, including random crop (or random resize crop for ImageNet-1K) and random horizontal flip (“weak aug”). Second, we use a 10-epoch warmup procedure for strong data augmentation. The warmup of Randaugment is implemented by progressively increasing the distortion magnitude from 1 to 9 (“warmup aug”). For CutMix and MixUp, we warm up by increasing the mixup probability from 0.5 to 1.0. As shown in Figure 3, “weak aug” provides the best accuracy. The “warmup aug” shows a slightly improved accuracy compared to fusing strong augmentation. Therefore, we provide a different result from [6] on the smaller dataset CIFAR-10, i.e., we show that weak augmentation is better than warmup augmentation. Even data augmentation with reduced amplitude is still difficult to learn at the beginning of adversarial training. Although strong augmentation is harmful to a normal training schedule, we show in Table VI that CutMix [95], MixUp [97], and Randaugment [21] increase the accuracy of adversarial training when training with a longer schedule, e.g., 800 epochs of fine-tuning. We conjecture that the data with strong augmentation is helpful but difficult for adversarial training to learn. Thus, it needs more epochs to learn meaningful representation. Loss and accuracy curves can be found in Appendix C. Due to the longer schedule, we use patch size 4 to reduce training time.
TABLE VII: Comparison between HSIC and $I_{\alpha}$. Models are pre-trained for 800 epochs. For fine-tuning, all models are adversarially trained with 10-step PGD for 50 epochs.
@@ 4.3 Data Augmentation is Harmful
MIMIR 0.001 HSIC [34] 69.63 43.17
MIMIR 0.001 $I_{\alpha}$ [94] 75.00 46.11
MIMIR 1e-05 HSIC [34] 76.30 47.60
MIMIR 1e-05 $I_{\alpha}$ [94] 75.53 46.75
MIMIR 1e-06 HSIC [34] 74.90 46.19
MIMIR 1e-06 $I_{\alpha}$ [94] 74.60 45.66
Adv-MAE (1-step) 0.0 - 74.69 46.28
Adv-MAE (10-step) 0.0 - 73.96 45.77
@@ 4.4 MI Measures
In Section 3.4, we provide lower and upper bound (Eq. (14)) of $I(x+\delta,z)$. According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the clean image $x$. This motivates us to directly embed $I(x+\delta,z)$ as a minimizing learning objective. Before achieving the goal, we must choose a suitable and efficient MI estimator and a regularizer ($\lambda$ in Eq. (8)). In this paper, we use $I_{\alpha}$ [94] and HSIC [54] as estimators. In Table VII, we show the performance of $I_{\alpha}$ [94] and HSIC [54] with different values of $\lambda$. According to the results, using HSIC as an estimator with $\lambda=1e-05$ shows the best performance, which is used for other experiments.

### v3

@@ 3.3 Design Details
where $\theta$ are the autoencoder parameters. After the autoencoder is trained, we discard the decoder and initialize a classification layer for the encoder to build a complete model. Finally, the complete model is fine-tuned by AT methods.
@@ 3.4 Theoretical Justification
Next, we provide theoretical justification showing that the MI between the adversarial example and its latent representation, i.e., $I(x+\delta,z)$, should be constrained. Let $F$ denote any classifier trained on natural samples with desirable prediction accuracy, which may suffer from adversarial attacks. We begin our analysis by first presenting Lemma 3.1.
Lemma 3.1.
Let $F(x+\delta)$ and $F(x_{re})$ denote, respectively, the predicted labels of adversarial sample $x+\delta$ and reconstructed sample $x_{re}$, we have:
@@
which implies that $F(x+\delta)$ is an indirect observation of $x+\delta$, whereas both $F(x_{re})$ and $x_{re}$ are indirect observations of $z$.
@@
Now, we define $p_{e}$ as the probability that the predicted label of $x+\delta$ by $F$ is not equal to that of $x_{re}$, i.e., $p_{e}=\mathbb{P}(F(x+\delta)\neq F(x_{re}))$. Intuitively, our autoencoder is trained to recover only natural sample $x$ without any interference from $\delta$. Hence, a relatively large value of $p_{e}$ is expected. In the following, we establish the connection between $p_{e}$ and $I(x+\delta,z)$ with both lower and upper bounds, showing that minimizing $I(x+\delta,z)$ also encourages a large value of $p_{e}$.
[…]
Architecture Params (M) FT FT Epoch Adv. Steps Source Natural Robust
ResNet-50 25.0 PGD 100 1 (Bai et al. 2021) 67.44 35.54
DeiT-S 22.1 PGD 100 1 (Bai et al. 2021) 66.62 36.56
DeiT-S 22.1 PGD 110 1 (Debenedetti, Sehwag, and Mittal 2023) 66.80 37.90
ViT-S 22.1 PGD 120 3 (Mao et al. 2022) 66.43 39.20
ViT-S 22.1 PGD 300 3 (Liu et al. 2024) 70.7 43.7
RobArch-S 26.1 PGD 110 3 (Peng et al. 2023) 70.17 44.14
ViT-S 22.1 APGD 300 2 (Singh, Croce, and Hein 2023) 69.22 44.04
ViT-S 22.1 APGD 300 2 Ours 71.00 46.10
ViT-B 86.6 ARD+PRM 10 5 (Mo et al. 2022) 69.10 34.62
ViT-B 86.6 PGD 120 3 (Mao et al. 2022) 70.64 43.04
RobArch-L 104 PGD 100 3 (Peng et al. 2023) 73.44 48.94
ViT-B 86.6 PGD 300 3 (Liu et al. 2024) 74.7 49.7
[…]
ViT-B 86.6 APGD 300 2 (Singh, Croce, and Hein 2023) 74.10 50.30
ViT-B 86.6 APGD 300 2 Ours 76.32 54.28
Proposition 3.2.
@@
where $|F(x+\delta)|$ is the total number of categories. [^2] For instance, for CIFAR-10, $|F(x+\delta)|=10$.
Proof.
@@
By applying Fano’s inequality (Fano 1949; Ocal et al. 2019), we obtain:
@@
Therefore, we obtain a lower bound of $I(x+\delta,z)$. If we use CIFAR-10 ($|F(x+\delta)|=10$) and assume the predicted labels $F(x+\delta)$ follow a uniform distribution, we can visualize the lower bound as a function of $p_{e}$ as shown in Figure 2, from which we observe an obvious monotonic inverse relationship between $I(x+\delta)$ and $p_{e}$ in the range $p_{e}\in[0,0.9]$. In fact, we can also obtain an upper bound, under the assumption that $I(F(x+\delta),x_{re})\approx I(x+\delta,z)$, i.e., there is no information loss in the two Markov chains in Eq. (8).
Proposition 3.3.
@@
Proof.
By the Hellman-Raviv inequality (Hellman and Raviv 1970; Brown 2009), we have:
@@
Similar to the lower bound, the upper bound also indicates $I(x+\delta,z)$ is inversely proportional to $p_{e}$ as shown in Figure 2. In fact, apart from the above-mentioned lower and upper bounds, there exists an alternative and intuitive way to understand the mechanism of minimizing $I(x+\delta,z)$. For simplicity, let us assume the natural data $x$ and adversarial perturbations $\delta$ are independent [^3] This assumption is mild for certain scenarios, such as when considering universal or image-agnostic perturbations (Moosavi-Dezfooli et al. 2017)., then:
@@
According to (Vincent et al. 2008), minimizing the expected reconstruction error between natural sample $x$ and corrupted input $x+\delta$ amounts to maximizing a lower bound of the mutual information $I(x,z)$, even though $z$ is a function of the corrupted input. Therefore, by minimizing $I(x+\delta,z)$, the network is forced to minimize $I(\delta,z)$ (since $I(x,z)$ is maximized). In other words, only the adversarial information about $\delta$ has been removed from $z$ when minimizing $I(x+\delta,z)$. This also explains the robustness of $z$.
4 Experiments
4.1 Experimental Setup
We evaluate MIMIR on three datasets: ImageNet-1K (Deng et al. 2009), Tiny-ImageNet (Le and Yang 2015), and CIFAR-10 (Krizhevsky, Hinton et al. 2009). Details of datasets are provided in Appendix A. Hyperparameters of the decoder are included in Appendix B.
Training Setup. We train models from scratch for all experiments. Following MAE (He et al. 2022), we do pre-training by MIMIR for 800 epochs. The number of warmup epochs is 40 for pre-training. We use AdamW (Loshchilov and Hutter 2019) as an optimizer for both pre-training and fine-tuning. The cosine decay as the learning rate scheduler is applied. At pre-training, MIMIR uses the 1-step PGD to generate adversarial examples for all three datasets. The perturbation budget is $\epsilon=8,\alpha=10$. At fine-tuning stage, we use the 10-step PGD AT with perturbation bound $\epsilon=8,\alpha=2$ for CIFAR-10 and Tiny-ImageNet. For ImageNet-1K, we use 1-step PGD with random initialization for better efficiency. The perturbation bound is $\epsilon=4$ for ImageNet-1K. We also use a 2-step APGD with a longer fine-tuning scheduler to compare with SOTA works in Table 1. Details on training hyperparameters are provided in Appendix C.
Evaluation Metrics. We use natural accuracy and robust accuracy as evaluation metrics. Natural accuracy refers to the accuracy of natural and unmodified inputs. The robust accuracy measures the accuracy under the AutoAttack (Francesco and Hein 2020). AutoAttack is an ensemble of diverse parameter-free attacks, including white-box and black-box attacks. In our experiments, we use standard version of AutoAttack that contains four attacks, including APGD-ce (Francesco and Hein 2020), APGD-t (Francesco and Hein 2020), FAB-t (Croce and Hein 2020), and Square (Andriushchenko et al. 2020). The perturbation budgets are $\epsilon=8$ for CIFAR-10 and Tiny-ImageNet, $\epsilon=4$ for ImageNet-1K.
4.2 Experimental Results
Comparison with SOTA Works. Table 1 compares MIMIR with previous works concerning adversarial robustness on ImageNet-1K ($l_{\infty},\epsilon=4/255$). We report the number of parameters, training epochs, steps in the inner maximization of AT, and clean and robust accuracy. The robust accuracy is evaluated by AutoAttack on the RobustBench validation set (5000 images). We divide the models into two groups: small ($\approx 22$M) and large ($\approx 86$M) models. In addition, (Singh, Croce, and Hein 2023) shows that the Convolutional stem (ConvStem) improves the robustness of ViTs. We also include an evaluation of ViTs+ConvStem in Table 2. ConvStem (Xiao et al. 2021) refers to replacing the patch embedding in ViTs with a convolutional block. MIMIR outperforms all SOTA works.
Table 2: Influence of the ConvStem block. The Standard pre-trained weights refer to pre-trained weights from timm. The configurations are deit_small_patch16_224 and vit_base_patch16_224.augreg_in1k. The weights are fine-tuned 300 epochs by 2-step APGD. †The result on ViT-S is taken from (Singh, Croce, and Hein 2023).
Architecture Pre-trained weights Natural Robust
ViT-S+ConvStem Standard 72.56 48.08†
MIMIR 72.72 48.44
ViT-B+ConvStem Standard 75.26 52.70
MIMIR 76.72 54.04
Ablation Study. Table 3 provides an ablation study to verify the design choices of MIMIR. The ablation uses 100 epochs of 1-step PGD (PGD1) AT as the baseline. Then, we apply end-to-end clean ImageNet-1K pre-training (weights available in timm library) as initialization of AT. After that, we replace the clean pre-training with MAE, adv MAE, and MIMIR step by step. The adv MAE refers to using adversarial examples but not using the MI $I(x+\delta,z)$ in loss. The pre-training scheduler is 800 epochs. We also use stronger adversarial fine-tuning for better performance, including 2-step PGD (PGD2), APGD (APGD2) FT, and a longer fine-tuning scheduler (300 epochs). Please note that a longer training scheduler does not guarantee better results (Mo et al. 2022). Our results indicate that MIMIR outperforms baselines and can be further improved under the long training scheduler.
In addition, MIMIR is also efficient and increases less than 5% of time consumption on ImageNet-1K compared to MAE, which is shown in Table 12 in the Appendix E.
@@ 4.2 Experimental Results
Architecture Training Recipe Natural Robust
ViT-S PGD1 FT w/o PT 66.02 31.40
clean PT + PGD1 FT 67.04 33.70
MAE PT + PGD1 FT 69.98 35.64
@@
MIMIR PT + PGD1 FT 71.02 37.22
MIMIR PT + PGD2 FT 70.78 38.16
MIMIR PT + APGD2 FT 68.78 42.86
100 $\rightarrow$ 300 epochs FT 71.00 46.10
Table 4: Comparison between HSIC (Gretton et al. 2005) and $I_{\alpha}$ (Yu, Yu, and Príncipe 2021) using ViT-T on CIFAR-10. Models are pre-trained 800 epochs and adversarially fine-tuned with 10-step PGD for 50 epochs.
@@
MIMIR 0.001 HSIC 69.63 43.17
MIMIR 0.001 $I_{\alpha}$ 75.00 46.11
MIMIR 1e-05 HSIC 76.30 47.60
MIMIR 1e-05 $I_{\alpha}$ 75.53 46.75
MIMIR 1e-06 HSIC 74.90 46.19
MIMIR 1e-06 $I_{\alpha}$ 74.60 45.66
Table 5: Comparison between different pre-training settings. All models are pre-trained for 800 epochs and then fine-tuned with 10-step PGD for 50 epochs using ViT-T on CIFAR-10.
Pre-train $\lambda$ Estimator Natural PGD
@@
MI Measure. In Section 3.4, we provide lower and upper bound (Eq. (11)) of $I(x+\delta,z)$. According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the natural image $x$. This motivates us to directly embed $I(x+\delta,z)$ as a minimizing learning objective. In this paper, we use $I_{\alpha}$ (Yu, Yu, and Príncipe 2021) and HSIC (Ma, Lewis, and Kleijn 2020) as estimators (detailed definitions in Appendix F). Table 4 demonstrates the performance of different values of $\lambda$. According to the results, we use HSIC with $\lambda=1e-05$ for all other experiments.